# HAFO: Humanoid Force-Adaptive Control for Intense External Force Interaction Environments

### v2

@@ 3.3 Upper-body motion generation based on imitation learning
The demonstration data for the upper-body policy come from the AMASS human-motion dataset after feasibility filtering. Following PHC [24], we retarget human poses to humanoid poses and apply filtering to suppress jitter and discontinuities. Instead of open-loop replay, we train a reinforcement-learning policy for the upper-body that can adapt to changing environments, improve whole-body coordination. The upper-body policy $\pi^{\mathrm{u}}:C^{{u}}\times S\mapsto A^{{u}}$ is trained in conjunction with the lower-body policy $\pi^{1}:C^{l}\times S\mapsto A^{l}$, and the state space of the upper-body strategy contains the state of the whole body. To improve the load capacity of the humanoid robot, we apply disturbances to the robot’s end-effectors in the form of random 3D spatial forces [6].
When the training of the upper-body strategy, samples are randomly taken from the offline dataset $M_{{t}}=\mathrm{q}_{1:T}^{\text{upper }}$ to keep the data variety and prevent overfitting to local patterns. To reduce the learning difficulty of the lower-body policy, we design a progressive imitation curriculum that guides the network gradually learns the upper-body reference trajectory, progressing from simple to complex and coarse to fine.
@@
Among them, $q_{0}(t)$ represents the actual joint angle vector at time t, $q_{\text{tar}}(t)$ is the target joint angle vector of the upper-body at that time $t$, and $\alpha_{i}\in[0,1]$ is the course gain coefficient, which is used to modulate the motion amplitude of the upper-body reference trajectory.
We adopt an asymmetric actor-critic training approach, where the critic network used both observable information $o_{t}$ and privileged information $o_{o}^{\text{priv}}=\left\{F_{\text{ee}},F_{\text{torso}},v_{\text{lin}}\right\}$ to achieve accurate value estimation, while the actor network relies solely on the observable information of the physical machine.
3.4 Impedance control
In complex and unpredictable environments, a precise and stable control system is essential for achieving task objectives. To enhance the robustness of humanoid robots under unknown external disturbances, this paper proposes a pulling-force modeling mechanism based on impedance control [43, 44, 45], explicitly incorporating the pulling force interaction process into the training loop. We extend the principle of impedance control to humanoid robots, aiming to regulate the robot’s body dynamics through a virtual spring-damper model, the corresponding dynamic model is given by
@@ 3.4 Impedance control
Where $K_{p}$ and $K_{d}$ represent the stiffness and damping of the virtual spring-damper system, $\left(\mathrm{x}_{\mathrm{des}},\dot{\mathrm{x}}_{\mathrm{des}}\right)$ denotes the desired position and velocity.
@@ 4.1 Force adaptive control under hand-induced disturbances
To evaluate the stability performance of the model under external disturbances, we apply forces of different directions and magnitudes to the robot’s end-effectors in the simulation environment, with the force magnitude incrementally increasing from 10N to 50N. During the force disturbance process, the robot executed random velocity commands (linear velocity range [-0.5, 0.5] m/s, angular velocity range [-0.5, 0.5] rad/s) and performed diverse upper-limb motion sequences sampled from a subset of the AMASS dataset. We test key indicators such as the upper-body motion tracking error and velocity tracking error. Motion tracking error is measured by calculating the difference between the actual joint angles and the target joint angles, and velocity tracking error is measured by calculating the difference between the actual velocity and the target velocity.
@@
The velocity tracking error $\mathrm{E}_{\text{error }}^{\text{vel }}(\mathrm{m}/\mathrm{s})$ and upper-body motion tracking error $\mathrm{E}_{\text{error }}^{\text{upper }}(\mathrm{rad})$ are computed as the average values over a complete iteration. To verify the effectiveness of the proposed Force-Adaptive Control strategy, we select two representative benchmarks:
@@
Table 1: Policy Tracking Performance Evaluation. We assessed the upper-body motion tracking performance and lower-body velocity tracking performance of different policies under three external force magnitudes applied to each hand (N-Force: 10 N, S-Force: 30 N, M-Force: 50 N). These metrics reflect the force adaptability of the policies under external disturbance, where ”/” indicates task failure.
@@
Furthermore, we conduct ablation experiments to quantify the contribution of each HAFO module. We sequentially ablate (i) the dual-agent architecture, retaining only the lower-body agent (w.o. DA), and (ii) the force-adaptation module (w.o. Force). As shown in Table 1, comparative experiments demonstrate that HAFO achieves the most accurate tracking performance in terms of velocity tracking and upper-body angle following. Upper-FIX-Lower-RL simplifies training by fixing the upper-body and relying exclusively on lower-body control for locomotion; however, it cannot guarantee system stability when the upper-body execute dynamic motions. Notably, the policy demonstrates a more pronounced performance advantage as the hand-applied external force increases. Furthermore, ablation results demonstrate that removing either module leads to performance degradation. The random upper-body motion introduced during training serves as an adversarial perturbation source, potentially compelling the lower-body policy to learn more generalized dynamic compensation mechanisms. Meanwhile, force curriculum learning effectively enhances the dual-agent’s robustness and adaptability under perturbations.
@@ 4.3 Force-Adaptive Control under Thrust Disturbance
Figure 3: Policy performance under varied external disturbances. (a) Performance under continuous constant-force disturbances applied from multiple directions. (b) Performance under one-second transient force disturbances from multiple directions.
To systematically evaluate policy robustness against thrust perturbations, we designed two disturbance rejection experiments: applying sustained constant forces and one-second transient forces along eight directions in the horizontal plane, progressively increasing force magnitude until loss of stability. As shown in Fig 3, HAFO achieved greater tolerance to external disturbances in most directions compared to HAFO w.o. DA, HAFO w.o. Force, and the baseline strategy, indicating that adversarial upper-body training and the force curriculum both enhance the robot’s force adaptation under thrust perturbations.
@@ 5 Real-world experiments
We deployed the HAFO strategy on physical Unitree G1 humanoid robot and conducted a series of hardware experiments. As shown in Fig.4, the HAFO controller accomplished multiple force interaction tasks, collectively validating our method’s technical performance and practical applicability on the physical robot.
Figure 4: HAFO enables force-adaptive control for humanoid robots across multiple scenarios. (a) Whole-body locomotion with sandbag loads. (b) Walking while carrying a box with one hand. (c) Being lifted onto an elevated platform and resuming locomotion. (d) Walking while carrying flower pots with both hands. (e) Policy deployment in a suspended state. (f) Performing dance movements. (g) Window cleaning while suspended by ropes.
Table 3: Policy Tracking Performance Evaluation in Real-world robot. We evaluate upper-body motion tracking error and lower-body velocity tracking error of different policies across difficulty hand-loading modes (1-kg load in one hand, 1-kg load in each hand).
@@ 5.1 Loco-manipulation under hand loads
To evaluate the robot’s motion performance under load, we conduct experiments with both single-hand and dual-hand loads. In the experiment, we record the velocity tracking error and upper-body motion tracking error respectively. As shown in Table 3, whether the robot carried the load with one hand or both, HAFO maintain the lowest velocity-tracking and upper-limb joint errors, outperforming the baseline methods.
Table 4: Policy performance evaluation across different task scenarios. We assessed motion smoothness and upper-limb tracking error for each task scenario.
@@ 5.2 Stable control under suspension
Traditional policy initialization schemes typically require robots to sequentially execute the series of procedures: foot-ground contact, posture adjustment, policy execution, and rope release. This process is relatively cumbersome, and sustained external disturbances from the rope readily leads to robot instability. This experiment validates the feasibility of suspended startup. As shown in Table 4 , the HAFO controller can initiate directly from a rope-suspended state and rapidly reach steady state, significantly simplifying the policy deployment process. Under the dynamic interference of rope tension, significant changes occurred in the robot’s locomotion states, including linear velocity, angular velocity, body posture, and foot-ground contact, which lead to leads to instability in the Baseline policy. In contrast, the HAFO policy leverages environmental feedback for adaptive compensation, robustly maintaining locomotion stability.
To verify the force adaptability in the rope suspension scenario, we simulate the typical suspended operation conditions, and systematically test the whole process of lifting, operating and landing. In the lifting phase, the robot is lifted from the ground, and its posture stability is monitored in real time. After entering the operation phase, the robot performs the teleoperation tasks in the suspended state. Finally, the robot achieves a stable bipedal stance following touchdown. We compare with the existing scheme, and finally only the HAFO completed the final experiment, the results are shown in Table 4. We show the application potential of HAFO in high-rise outdoor dangerous scenes, and provide new ideas and methods for the practical application of robots in the future.
@@ 6 Conclusion
In this paper, we propose HAFO, a force-adaptive control framework for humanoid robots. HAFO employs a dual-agent control strategy and explicitly introduces various dynamic disturbances, achieving stable motion control and precise motion imitation. During the model training, a spring-damper system is established, and fine-grained tension control of the pulling force is achieved through manipulation of virtual springs. At the same time, HAFO employs asymmetric Actor-Critic architecture wherein the critic-network is granted full access to privileged state variables, thereby furnishing a more accurate approximation of state-action value. Experimental results demonstrate that the proposed strategy enables the humanoid robot to exhibit strong adaptability under challenging conditions such as heavy loading, thrust disturbance and rope suspension, providing robust support for solving substantial force interaction tasks in real-world applications.

### v3

@@ 3.3 Upper-body motion generation based on imitation learning
The demonstration data for the upper-body policy are derived from the AMASS human-motion dataset [43] after feasibility filtering. Following PHC [24], human poses are retargeted to humanoid poses, and filtering is applied to suppress jitter and discontinuities. Instead of using open-loop replay, we train a reinforcement learning policy for the upper body that adapts to changing environments, avoids self-collisions, and improves whole-body coordination. The upper-body policy $\pi^{\mathrm{u}}:C^{{u}}\times S\mapsto A^{{u}}$ is trained in conjunction with the lower-body policy $\pi^{l}:C^{l}\times S\mapsto A^{l}$, and the state space of the upper-body strategy includes the state of the entire body. To enhance the load capacity of the humanoid robots, disturbances are applied to the robot’s end-effectors in the form of random 3D spatial forces [6].
During the training of the upper-body strategy, samples are randomly selected from the offline dataset $M_{{t}}=\mathrm{q}_{1:T}^{\text{upper }}$ to maintain data variety and prevent overfitting to local patterns. To alleviate the learning difficulty of the lower-body policy, we design a progressive imitation curriculum that guides the network to gradually learn the upper-body reference trajectory, progressing from simple to complex and coarse to fine.
@@
Among them, $q_{0}(t)$ represents the actual joint angle vector at time t, $q_{\text{tar}}(t)$ is the target joint angle vector of the upper-body at that time $t$, and $\alpha_{i}\in[0,1]$ is the course gain coefficient, which modulates the motion amplitude of the upper-body reference trajectory.
We adopt an asymmetric actor-critic training approach, where the critic networks use both observable information $o_{t}$ and privileged information $o_{o}^{\text{priv}}=\left\{F_{\text{ee}},F_{\text{torso}},v_{\text{lin}}\right\}$ to achieve accurate value estimation, while the actor networks rely solely on the observable information of the physical machine.
3.4 Virtual spring-damper system
In complex and unpredictable environments, a precise and stable control system is crucial for achieving task objectives. To improve the robustness of humanoid robots in the presence of unknown external disturbances, this paper introduces a pulling-force modeling mechanism based on spring-damper system [44, 45, 46], explicitly incorporating the pulling force interaction process into the training loop, aiming to regulate the robot’s body dynamics through a virtual spring-damper model, the corresponding dynamic model is given by
@@ 3.4 Virtual spring-damper system
Where $K_{p}$ and $K_{d}$ represent the stiffness and damping of the virtual spring-damper system, $\left(\mathrm{x}_{\mathrm{des}},\dot{\mathrm{x}}_{\mathrm{des}}\right)$ denotes the desired position and velocity, and $f_{spring}$ represents the spring-damper force.
@@ 4.1 Force adaptive control under hand-induced disturbances
To evaluate the stability performance of the model under external disturbances, we apply forces of different directions and magnitudes to the robot’s end-effectors in the simulation environment, with the force magnitude incrementally increasing from 10N to 50N. During the force disturbance process, the robot executes random velocity commands (linear velocity range: [-0.5, 0.5] m/s, angular velocity range: [-0.5, 0.5] rad/s) while performing diverse upper-limb motion sequences sampled from a subset of the AMASS dataset [43]. We test key indicators such as the upper-body motion tracking error and velocity tracking error. Motion tracking error is quantified by calculating the difference between the actual joint angles and the target joint angles, and velocity tracking error is quantified by calculating the difference between the actual velocity and the target velocity.
@@
The velocity tracking error $\mathrm{E}_{\text{error }}^{\text{vel }}(\mathrm{m}/\mathrm{s})$ and upper-body motion tracking error $\mathrm{E}_{\text{error }}^{\text{upper }}(\mathrm{rad})$ are computed as the average values over a complete iteration. To evaluate the effectiveness of the proposed Force-Adaptive Control strategy, two representative benchmarks are selected:
@@
Table 1: Policy Tracking Performance Evaluation. We evaluated the upper-body motion tracking performance and lower-body velocity tracking performance of different policies under three external force magnitudes applied to each hand (N-Force: 10 N, S-Force: 30 N, M-Force: 50 N). These metrics reflect the force adaptability of the policies under external disturbance, where ”/” indicates task failure.
@@
Furthermore, ablation experiments are conducted to quantify the contribution of each HAFO module. We sequentially ablate (i) the dual-agent architecture, retaining only the lower-body agent (w.o. DA), and (ii) the force-adaptative module (w.o. Force). As shown in Table 1, comparative experiments demonstrate that HAFO achieves superior tracking performance, both in terms of velocity tracking and upper-body angle following. The Upper-FIX-Lower-RL policy simplifies training by fixing the upper-body and relying exclusively on lower-body control for locomotion. However, it fails to guarantee system stability when the upper-body executes dynamic motions in real-world scenarios. Notably, HAFO demonstrates a more pronounced performance advantage as the hand-applied external force increases. Furthermore, ablation results demonstrate that removing either module leads to performance degradation. The random upper-body motion introduced during training serves as an adversarial perturbation source, potentially compelling the lower-body policy to learn more generalized dynamic compensation mechanisms. Meanwhile, force-interaction curriculum learning effectively enhances the dual-agent’s robustness and adaptability under perturbations.
@@ 4.3 Force-Adaptive Control under Thrust Disturbance
Figure 3: Policy performance under various external thrust disturbances. (a) Performance under continuous constant-force disturbances applied from multiple directions. (b) Performance under one-second transient force disturbances from multiple directions.
To systematically assess policy robustness against thrust perturbations, we conduct two disturbance rejection experiments: applying sustained constant forces and one-second transient forces along eight directions in the horizontal plane, progressively increasing force magnitude until loss of stability. As shown in Figure 3, HAFO demonstrates greater tolerance to external disturbances in most directions compared to HAFO w.o. DA, HAFO w.o. Force, and the baseline strategy, indicating that adversarial upper-body training and the force curriculum both improve the robot’s force adaptation under thrust perturbations.
@@ 5 Real-world experiments
We deployed the HAFO strategy on physical Unitree G1 humanoid robot and conduct a series of hardware experiments. As shown in Figure 4, the HAFO controller completes multiple force interaction tasks, collectively validating the technical performance and practical applicability of our method on the physical robot.
Figure 4: HAFO enables force-adaptive control for humanoid robots across multiple scenarios. (a) Whole-body control with sandbag loads. (b) Locomotion while carrying a box with one hand. (c) Being lifted onto an elevated platform and resuming locomotion. (d) Locomotion while carrying flower pots with both hands. (e) Policy deployment in a suspended state. (f) Performing dance movements. (g) Window cleaning while suspended by ropes.
Table 3: Policy Tracking Performance Evaluation on the Real-world robot. We evaluate upper-body motion tracking error and lower-body velocity tracking error of different policies across difficulty hand-loading modes (1 kg load in one hand, 1 kg load in each hand).
@@ 5.1 Loco-manipulation under hand loads
To evaluate the robot’s motion performance under load, we conduct experiments with both single-hand and dual-hand loads. In the experiment, we record the velocity tracking error and upper-body motion tracking error respectively. As shown in Table 3, whether the robot carries the load with one hand or both, HAFO maintains the lowest velocity-tracking and upper-limb joint errors, outperforming the baseline methods.
Table 4: Policy performance evaluation across different task scenarios. We assess motion smoothness and upper-body tracking error for each task scenario.
@@ 5.2 Stable control under suspension
Traditional policy initialization schemes typically require robots to sequentially execute the series of procedures: foot-ground contact, posture adjustment, policy execution, and rope release. This process is relatively cumbersome, and sustained external disturbances from the rope readily leads to robot instability. This experiment validates the feasibility of suspended startup. The HAFO controller can initiate directly from a rope-suspended state and rapidly reach steady state, significantly simplifying the policy deployment process. Humanoid robots are typically controlled based on high-level signals such as target velocity and foot contact commands. Under the dynamic interference of rope tension, significant changes occur in the robot’s locomotion states, including linear velocity, angular velocity, body posture, and foot-ground contact. The baseline strategy misinterprets the normal swinging of the rope as robot instability, triggering emergency actions that cause violent oscillations. In contrast, the HAFO policy leverages environmental feedback for adaptive compensation, robustly maintaining locomotion stability.
To verify the force adaptability in the rope suspension scenario, we simulate typical suspended operation conditions and systematically test the whole process of lifting, operating and landing. During the lifting phase, the robot is lifted from the ground, and its posture stability is monitored in real time. After entering the operation phase, the robot performs the teleoperation tasks in the suspended state. Finally, the robot achieves a stable bipedal stance following touchdown. We compare with the existing scheme, and finally only the HAFO completed the final experiment. The results are shown in Table 4, we show the application potential of HAFO in high-rise outdoor dangerous scenes, and provide new ideas and methods for the practical application of robots in the future.
@@ 6 Conclusion
In this paper, we propose HAFO, a force-adaptive control framework for humanoid robots. HAFO employs a dual-agent control strategy and explicitly introduces various dynamic disturbances, achieving stable motion control and precise motion imitation. During the model training, a spring-damper system is established, and fine-grained tension control of the pulling force is achieved through manipulation of virtual springs. Additionally, HAFO employs asymmetric Actor-Critic architecture wherein the critic-network is granted full access to privileged state variables, thereby furnishing a more accurate approximation of state-action value. Experimental results demonstrate that the proposed strategy enables the humanoid robots to exhibit strong adaptability under challenging conditions such as heavy load, thrust disturbance and rope suspension, providing robust support for solving substantial force interaction tasks in real-world applications.
Appendix A RL training
We use Proximal Policy Optimization (PPO) [41] as the base algorithm, and Table A.1 presents the detailed training hyperparameters.
Table 1: RL Controller Hyperparameters.
Hyperparameter Value
Clip param 0.2
Discount factor($\gamma$) 0.99
Value loss coef 1.0
Entropy coef 0.01
Actor learning rate 0.001
Critic learning rate 0.001
Weight decay 0.01
Max grad norm 1.0
Desired KL 0.01
Num steps per env 24
MLP size [512, 256, 128]
Appendix B State and Action Space
This section provides a detailed description of the state and action spaces for the dual-agent policy. The state space for both the upper and lower body strategies is identical, consisting of the whole-body state set. The dimensions of the action space for the upper and lower body strategies correspond to the number of degrees of freedom (DOF) in the humanoid robot’s upper and lower bodies, respectively. Specific details are shown in Table B.2.
Table 2: State and Action space description.
Term Lower dim Upper dim Whole dim
Base ang vel 3 3 3
Projected gravity 3 3 3
Commands 5 (Vel+Stand) 14 (Motion angle) 19
Dof pos 29 29 29
Dof vel 29 29 29
Actions 12 (Lower) 15 (Upper) 27 (Whole)
Total dim 81 93 110
We adopt an asymmetric Actor-Critic architecture for reinforcement learning training. In this architecture, the Critic network has access to privileged information, including the base linear velocity and external force values, which allows it to more accurately evaluate the value of the current state and provide more precise feedback signals to the Actor network.
Appendix C Reward Design
This paper adopts a reward function design similar to that in [2, 6, 19], while making targeted additions and modifications to certain reward components to better accommodate the needs of full-body control tasks and force adaptation control. These adjustments aim to enhance both the training effectiveness and the robustness of the strategy. The specific reward components are detailed in Table C.3.
Table 3: Definition of the additional reward function
Term Expression Weight
Regularization Reward
Orientation $\|\mathbf{g}_{xy}^{\text{root}}\|_{2}^{2}$ -4.0
Torso Orientation $\|\mathbf{g}_{xy}^{\text{torso}}\|_{2}^{2}$ -7.0
Lower-Body Action Rate $\|a_{t}^{\text{lower}}-a_{t-1}^{\text{lower}}\|^{2}$ -0.2
Feet Orientation $\|\mathbf{g}_{xy}^{\text{feet}}\|_{2}^{2}$ -4.0
Termination $\mathbb{I}_{\text{termination}}$ -350
Feet Parallel $\operatorname{Var}(D)$ -2.0
Lower-Body Stand Still $\|q_{t}^{\text{upper}}-q_{t}^{\text{default}}\|\times\mathbb{I}(\|F_{\text{spring}}\|)>0$ -2.2
Base Height $\|\mathbf{h}-\mathbf{h}_{\text{target}}\|_{2}^{2}\times\mathbb{I}(\|F_{\text{spring}}\|)=0$ -10
Additional Torques $\|\boldsymbol{\tau}\|_{2}^{2}\times\mathbb{I}(\|F_{\text{spring}}\|)>0$ -0.0001
Additional Dof Vel $\|\dot{\mathbf{q}}\|_{2}^{2}\times\mathbb{I}(\|F_{\text{spring}}\|)>0$ -0.008
Additional Dof Acc $\|\ddot{\mathbf{q}}\|_{2}^{2}\times\mathbb{I}(\|F_{\text{spring}}\|)>0$ -0.000011
Additional Action Rate $\|a_{t}-a_{t-1}\|_{2}^{2}\times\mathbb{I}(\|F_{\text{spring}}\|)>0$ -0.01
Horizontal Angular Velocity $\|\boldsymbol{\omega}_{xy}\|_{2}^{2}\times\mathbb{I}(\|F_{\text{spring}}\|)=0$ -1.0
Task Reward
Linear X Velocity $\exp[-\|v_{t}^{x*}-v_{t}^{x}\|^{2}/0.25]\times\mathbb{I}(\|F_{\text{spring}}\|)=0$ 2.0
Linear Y Velocity $\exp[-\|v_{t}^{y*}-v_{t}^{y}\|^{2}/0.25]\times\mathbb{I}(\|F_{\text{spring}}\|)=0$ 2.0
Angular Yaw Velocity $\exp[-\|v_{t}^{\text{ang}*}-v_{t}^{\text{ang}}\|^{2}/0.25]\times\mathbb{I}(\|F_{\text{spring}}\|)=0$ 6.0
Upper-Body Dofs Position $\exp[-\|q_{t}^{\text{upper}*}-q_{t}^{\text{upper}}\|^{2}/0.25]$ 4.0
Lower-Body Stand Still $\exp[-\|q_{t}^{\text{lower}}-q_{t}^{\text{default}}\|^{2}/0.25]\times\mathbb{I}(\|F_{\text{spring}}\|)>0$ 3.0
Appendix D Domain randomization
To enhance the robustness of the policy and reduce the performance gap between the simulation and real environments, we applied domain randomization to key parameters in the simulation. This approach allows the policy better adapt to unknown factors when transferred to the real environment. The specific domain randomization parameters and their ranges are listed in Table D.4.
Table 4: Domain randomization parameters and their ranges.
Term Value
Ground friction $\mathcal{U}(0.25,1.25)$
Kp and Kd $\mathcal{U}(0.9,1.1)$
Base mass $\mathcal{U}(-1,3)\,\mathrm{kg}$
Link mass $\mathcal{U}(0.9,1.2)\times\text{default}\,\mathrm{kg}$
Base CoM $\mathcal{U}(-0.01,0.01)\,\mathrm{cm}$
Control delay $\mathcal{U}(0,20)\,\mathrm{ms}$
Appendix E Extend HAFO to Large-size Humanoid Robots
As shown in Figure E.1, We further extend the HAFO strategy to the full-scale humanoid robot H1-2 (height: 1.78 m, weight: 70 kg). Due to its larger mass and limited actuators, this robot presents greater challenges in terms of control. However, with minimal parameter adjustments, stable control is achieved within the simulation environment, thereby validating the strategy’s strong generalization and scalability.
Figure 1: Multi-scenario validation of the HAFO policy on the unitree H1-2 humanoid robot. (a) Robot upper-body swing. (b) Punching and maintaining stable standing. (c) Stable locomotion with a 100 N external force applied to each hand. (d) Stable control in the suspended state. (e) Rapid robot adjustment after external impact.